# RoBridge: A Hierarchical Architecture Bridging Cognition and Execution for General Robotic Manipulation

### v2

@@ Abstract
Operating robots in open-ended scenarios with diverse tasks is a crucial research and application direction in robotics. While recent progress in natural language processing and large multimodal models has enhanced robots’ ability to understand complex instructions, robot manipulation still faces the procedural skill dilemma and the declarative skill dilemma in open environments. Existing methods often compromise cognitive and executive capabilities. To address these challenges, in this paper, we propose RoBridge, a hierarchical intelligent architecture for general robotic manipulation. It consists of a high-level cognitive planner (HCP) based on a large-scale pre-trained vision-language model (VLM), an invariant operable representation (IOR) serving as a symbolic bridge, and a generalist embodied agent (GEA). RoBridge maintains the declarative skill of VLM and unleashes the procedural skill of reinforcement learning, effectively bridging the gap between cognition and execution. RoBridge demonstrates significant performance improvements over existing baselines, achieving a 75% success rate on new tasks and an 83% average success rate in sim-to-real generalization using only five real-world data samples per task. This work represents a significant step towards integrating cognitive reasoning with physical execution in robotic systems, offering a new paradigm for general robotic manipulation.
@@ 1 Introduction
In this paper, we introduce RoBridge—a hierarchical intelligent architecture aimed at general robotic manipulation. As shown in Figure 1, our method maintains the declarative skill of VLM, and releases the procedural skill of reinforcement learning. This architecture innovatively constructs an intelligent system comprising three core elements: (1) a high-level cognitive planner (HCP) based on a large-scale pre-trained vision-language model; (2) an invariant operable representation (IOR) with physical intuition and environmental invariance, serving as a symbolic bridge between abstract cognition and embodied execution; (3) a generalist embodied agent (GEA) that achieves precise action execution through interaction with the physical world. By generating IOR through the causal reasoning capabilities of the HCP and transforming abstract symbols into operations in physical space via the GEA, RoBridge effectively resolves the challenge of the disconnection between cognition and execution. Our main contributions are as follows:
@@
  • Generalist Embodied Agent: We design a Generalist Embodied Agent (GEA) capable of converting invariant operable representation (IOR) into concrete execution actions, maintaining superior performance under various interference conditions.
@@ 3 Method
Figure 2: RoBridge overview. RoBridge adopts a three-layer architecture, consisting of a high-level cognitive planner (HCP), an invariant operable representation (IOR), and a generalist embodied agent (GEA). For example, for the instruction “Put the blocks into the corresponding shaped slots”, HCP will first plan and split the task into multiple primitive actions. Then, combined with the APIs composed of the foundation model, it will give IOR, which mainly includes the masked depth of the first perspective, the mask of the third perspective, the type of action, and the constraints. IOR is updated by HCP at a low frequency and track-anything updates the mask at a high frequency. IOR is used as the input of GEA, and GEA performs specific actions until the task is completed.
@@ 3.1 Framework
The framework of RoBridge is illustrated in Figure 2. It mainly contains three core components: a high-level cognitive planner, an invariant operation representation and a generalist embodied agent. First, the high-level cognitive planner decomposes tasks into primitive actions based on observation and instruction. Then, for each action, the HCP integrates foundation models to generate an invariant operation representation. Finally, the generalist embodied agent performs operations utilizing this representation, with the execution process being governed by closed-loop control. Detailed descriptions follow.
@@
Generalist Embodied Agent (GEA). At each time step, a new $\mathcal{R}^{t}_{i}$ is generated. The specific update process will be described in the next closed-loop control section. For each $\mathcal{R}^{t}_{i}$, we need to map it into robotic movement $\mathbf{a}_{t}$ to proceed with the primitive action. In other words, we need to learn a policy $\pi\left(\mathcal{R}^{t}_{i}\right)\mapsto\mathbf{a}_{t}$. Following PSL [12], the primitive action “reach” typically involves moving the robot’s end-effector to a well-defined target position. Unlike tasks such as grasping or placing, which involve complex object interactions and decision-making, “reach” can be efficiently addressed through motion planning. For other primitive actions, we integrate reinforcement learning and imitation learning to train a generalist embodied agent capable of performing actions effectively, while ensuring robustness and consistent performance under varying input conditions. Detailed descriptions are shown in Section 3.2.
@@
3.2 Generalist Embodied Agent (GEA) Training
Figure 3: Generalist Embodied Agent(GEA) Training. The left figure illustrates the domain randomization methods, with the first column showing the original images. The first row employs changes in robotic arm pose, object shape variations, and camera offsets, while the second row uses pixel offsets, depth distortion, hollowing, etc. The domain randomization in the first row is used during expert training, whereas all domain randomization methods are applied during GEA training. In the top right corner, it represents that we train an expert for each task, and the trained experts provide training data for GEA. The bottom right corner shows the training process of GEA, where it takes IOR as input and is trained using DAgger. Data from failed executions is corrected by the experts.
@@ 3.2 Generalist Embodied Agent (GEA) Training
IL Training. Next, we start training a generalist embodied agent $\pi_{g}$. We use expert agents to generate high-quality data for each task and extract generalized interaction representation $\mathcal{R}$ from the data as the input of $\pi_{g}$. In addition to the domain randomization employed during experts agent training, we further incorporated several domain randomization techniques, including depth distortion, dilation, random shifting, and the modification of masked information through addition and deletion. Such strategies are intended to improve the generalization capabilities and robustness of the trained agent across varying environmental conditions. To simulate the noisy depth sensors in the real world, we employ a series of augmentation techniques [13] when processing depth images. Specifically, we utilize depth warping, which applies Gaussian shifts to simulate variations in perspective and sensor noise. Additionally, we incorporate Gaussian blurring to mimic the effects of sensor blur and focus issues, which are common in real-world scenarios. Furthermore, we introduce random masking to create artificial holes in the depth maps, simulating occlusions and missing data that frequently occur in practical applications. Due to the potential inaccuracies and incomplete coverage inherent in the model-generated masks, we employ techniques such as random offsets and random cropping. These methods are applied to mitigate the risk of the agent becoming overly reliant on the masks.
@@ 4 Experiments
We first introduce the experimental setup 4.1 along with the benchmarks and baselines 4.2. We present the main results C and demonstrate performance on unseen tasks 4.4. Additionally, we conduct ablation studies 4.5 and analyze the causes of failures 4.6.
@@ 4.2 Benchmark and Baselines
Baselines. DrQ-v2 [54] is a state-of-the-art reinforcement learning method. We adapted a multi-task DrQ-v2 using RGB inputs, robot proprioception, and one-hot encoded tasks. SayCan [2] uses a large language model (LLM) for skill planning. We used DrQ-v2 models trained on separate tasks as SayCan’s skill library. PSL [12] organizes skills as actions, similar to SayCan. ManipGen [13] extends PSL with DAgger and domain randomization. We expanded PSL and ManipGen’s skill libraries for our tasks. ReKep [25] employs keypoint representations for task planning. Due to its initial suboptimal performance, we enhanced prompts with additional task specifications and constraints. We also compared popular end-to-end methods, including RDT [33], $\pi_{0}$ [5], and its autoregressive version, $\pi_{0}$-fast. For real-world testing of RDT, $\pi_{0}$, and our RoBridge, we collected 5 demonstrations per task for fine-tuning.
@@ 4.5 Ablation Study
We conducted ablation experiments on the components of the IOR and the training of the GEA to understand their significance. [13] used the original depth instead of the 1st masked depth, and [61] used features obtained from DINOv2 [39] as inputs instead of the mask in 3rd view. We conducted ablations on our modules using these two methods. As shown in the second and third rows of Table 5, the original depth might be easier for the model to learn due to the presence of more details, resulting in a higher success rate on the MT50 tasks. However, the success rate drops sharply when there are changes in camera pose. The performance of DINOv2 was not satisfactory, which we think is due to its pre-trained features not being fully suitable for robotics. The sharp decline in performance for both methods under unseen camera poses indicates that these representations lack good invariance, making it difficult to achieve generalization. We also conducted ablation studies on the methods employed in GEA training, as shown in the last three rows of Table 5. Each method was found to be critically important in enhancing the training effectiveness of the GEA.
4.6 Failure Case Analysis
@@ 5 Conclusion
In this paper, we introduce RoBridge, a novel hierarchical intelligent architecture designed to enhance robotic manipulation by bridging the gap between high-level cognitive planning and low-level physical execution. The architecture integrates a high-level cognitive planner, an invariant operable representation, and a generalist embodied agent, demonstrating significant advancements in task generalization and execution robustness. Through extensive experiments, RoBridge has shown superior performance and strong zero-shot generalization capabilities in unknown environments and novel tasks.
6 Acknowledgements
This work is supported by National Key Research and Development Program of China(2024YFE0203100), National Natural Science Foundation of China (NSFC) under Grants No.62476293, Shenzhen Science and Technology Program No.GJHZ20220913142600001, Nansha Key R&D Program under Grant No.2022ZD014, and General Embodied AI Center of Sun Yat-sen University, The Major Key Project of PCL (No. PCL2024A04, No. PCL2024AS204).
@@ B.2 Training Details
In our experiments on MetaWorld, we conducted training for 1M steps, with failure data sampling occurring every 100k steps. During real-world experiments, we fine-tuned the model for 2k steps. We will resize the image to 168$\times$168 and feed it to GEA for faster speed.

### v3

@@ Abstract
Operating robots in open-ended scenarios with diverse tasks is a crucial research and application direction in robotics. While recent progress in natural language processing and large multimodal models has enhanced robots’ ability to understand complex instructions, robot manipulation still faces the procedural skill dilemma and the declarative skill dilemma in open environments. Existing methods often compromise cognitive and executive capabilities. To address these challenges, in this paper, we propose RoBridge, a hierarchical intelligent architecture for general robotic manipulation. It consists of a high-level cognitive planner (HCP) based on a large-scale pre-trained vision-language model (VLM), an invariant operable representation (IOR) serving as a symbolic bridge, and a guided embodied agent (GEA). RoBridge maintains the declarative skill of VLM and unleashes the procedural skill of reinforcement learning, effectively bridging the gap between cognition and execution. RoBridge demonstrates significant performance improvements over existing baselines, achieving a 75% success rate on new tasks and an 83% average success rate in sim-to-real generalization using only five real-world data samples per task. This work represents a significant step towards integrating cognitive reasoning with physical execution in robotic systems, offering a new paradigm for general robotic manipulation.
@@ 1 Introduction
In this paper, we introduce RoBridge—a hierarchical intelligent architecture aimed at general robotic manipulation. As shown in Figure 1, our method maintains the declarative skill of VLM, and releases the procedural skill of reinforcement learning. This architecture innovatively constructs an intelligent system comprising three core elements: (1) a high-level cognitive planner (HCP) based on a large-scale pre-trained vision-language model; (2) an invariant operable representation (IOR) with physical intuition and environmental invariance, serving as a symbolic bridge between abstract cognition and embodied execution; (3) a guided embodied agent (GEA) that achieves precise action execution through interaction with the physical world. By generating IOR through the causal reasoning capabilities of the HCP and transforming abstract symbols into operations in physical space via the GEA, RoBridge effectively resolves the challenge of the disconnection between cognition and execution. While prior works have partially touched on individual ingredients, RoBridge is, to our knowledge, the first holistic system that synergistically unites the declarative power of large vision–language models with the procedural proficiency of reinforcement learning via the invariant operable representation (IOR) without mutual restriction. Our main contributions are as follows:
@@
  • Guided Embodied Agent: We design a Guided Embodied Agent (GEA) capable of converting invariant operable representation (IOR) into concrete execution actions, maintaining superior performance under various interference conditions.
@@ 3 Method
Figure 2: RoBridge overview. RoBridge adopts a three-layer architecture, consisting of a high-level cognitive planner (HCP), an invariant operable representation (IOR), and a guided embodied agent (GEA). For example, for the instruction “Put the blocks into the corresponding shaped slots”, HCP will first plan and split the task into multiple primitive actions. Then, combined with the APIs composed of the foundation model, it will give IOR, which mainly includes the masked depth of the first perspective, the mask of the third perspective, the type of action, and the constraints. IOR is updated by HCP at a low frequency and track-anything updates the mask at a high frequency. IOR is used as the input of GEA, and GEA performs specific actions until the task is completed.
@@ 3.1 Framework
The framework of RoBridge is illustrated in Figure 2. It mainly contains three core components: a high-level cognitive planner, an invariant operation representation and a guided embodied agent. First, the high-level cognitive planner decomposes tasks into primitive actions based on observation and instruction. Then, for each action, the HCP integrates foundation models to generate an invariant operation representation. Finally, the guided embodied agent performs operations utilizing this representation, with the execution process being governed by closed-loop control. Detailed descriptions follow.
@@
Guided Embodied Agent (GEA). At each time step, a new $\mathcal{R}^{t}_{i}$ is generated. The specific update process will be described in the next closed-loop control section. For each $\mathcal{R}^{t}_{i}$, we need to map it into robotic movement $\mathbf{a}_{t}$ to proceed with the primitive action. In other words, we need to learn a policy $\pi\left(\mathcal{R}^{t}_{i}\right)\mapsto\mathbf{a}_{t}$. Following PSL [12], the primitive action “reach” typically involves moving the robot’s end-effector to a well-defined target position. Unlike tasks such as grasping or placing, which involve complex object interactions and decision-making, “reach” can be efficiently addressed through motion planning. For other primitive actions, we integrate reinforcement learning and imitation learning to train a guided embodied agent capable of performing actions effectively, while ensuring robustness and consistent performance under varying input conditions. Detailed descriptions are shown in Section 3.2.
@@
3.2 Guided Embodied Agent (GEA) Training
Figure 3: Guided Embodied Agent(GEA) Training. The left figure illustrates the domain randomization methods, with the first column showing the original images. The first row employs changes in robotic arm pose, object shape variations, and camera offsets, while the second row uses pixel offsets, depth distortion, hollowing, etc. The domain randomization in the first row is used during expert training, whereas all domain randomization methods are applied during GEA training. In the top right corner, it represents that we train an expert for each task, and the trained experts provide training data for GEA. The bottom right corner shows the training process of GEA, where it takes IOR as input and is trained using DAgger. Data from failed executions is corrected by the experts.
@@ 3.2 Guided Embodied Agent (GEA) Training
IL Training. Next, we start training a guided embodied agent $\pi_{g}$. We use expert agents to generate high-quality data for each task and extract generalized interaction representation $\mathcal{R}$ from the data as the input of $\pi_{g}$. In addition to the domain randomization employed during experts agent training, we further incorporated several domain randomization techniques, including depth distortion, dilation, random shifting, and the modification of masked information through addition and deletion. Such strategies are intended to improve the generalization capabilities and robustness of the trained agent across varying environmental conditions. To simulate the noisy depth sensors in the real world, we employ a series of augmentation techniques [13] when processing depth images. Specifically, we utilize depth warping, which applies Gaussian shifts to simulate variations in perspective and sensor noise. Additionally, we incorporate Gaussian blurring to mimic the effects of sensor blur and focus issues, which are common in real-world scenarios. Furthermore, we introduce random masking to create artificial holes in the depth maps, simulating occlusions and missing data that frequently occur in practical applications. Due to the potential inaccuracies and incomplete coverage inherent in the model-generated masks, we employ techniques such as random offsets and random cropping. These methods are applied to mitigate the risk of the agent becoming overly reliant on the masks.
@@ 4 Experiments
We first introduce the experimental setup 4.1 along with the benchmarks and baselines 4.2. We present the main results C and demonstrate performance on unseen tasks 4.4. Additionally, we conduct ablation studies 4.5.
@@ 4.2 Benchmark and Baselines
Baselines. DrQ-v2 [54] is a state-of-the-art reinforcement learning method. We adapted a multi-task DrQ-v2 using RGB inputs, robot proprioception, and one-hot encoded tasks. SayCan [2] uses a large language model (LLM) for skill planning. We used DrQ-v2 models trained on separate tasks as SayCan’s skill library. PSL [12] organizes skills as actions, similar to SayCan. ManipGen [13] extends PSL with DAgger and domain randomization. We expanded PSL and ManipGen’s skill libraries for our tasks. RAM [30] is a retrieval-based, zero-shot robotic manipulation framework that transfers 2D affordances from diverse out-of-domain data to 3D executable actions. ReKep [25] employs keypoint representations for task planning. Due to its initial suboptimal performance, we enhanced prompts with additional task specifications and constraints. We also compared popular end-to-end methods, including RDT [33], $\pi_{0}$ [5], and its autoregressive version, $\pi_{0}$-fast. For real-world testing of RDT, $\pi_{0}$, and our RoBridge, we collected 5 demonstrations per task for fine-tuning.
@@ 4.3 Results Analysis
RoBridge w/ keypoints 80.2 62.2 55.4 58.4 43.2 59.88
RoBridge w/ language-only 84.8 34.4 26.8 30.2 20.2 39.28
RoBridge w/ smaller VLM 65.4 62.8 60.4 66.2 64.4 63.84
@@ 4.5 Ablation Study
We conducted ablation experiments on the components of the IOR and the training of the GEA to understand their significance. [13] used the original depth instead of the 1st masked depth, and [62] used features obtained from DINOv2 [39] as inputs instead of the mask in 3rd view. We conducted ablations on our modules using these two methods. As shown in the second and third rows of Table 5, the original depth might be easier for the model to learn due to the presence of more details, resulting in a higher success rate on the MT50 tasks. However, the success rate drops sharply when there are changes in camera pose. The performance of DINOv2 was not satisfactory, which we think is due to its pre-trained features not being fully suitable for robotics. The sharp decline in performance for both methods under unseen camera poses indicates that these representations lack good invariance, making it difficult to achieve generalization. The comparison with using keypoints and language shows that RoBridge’s success is not only due to richer supervision but also because the IOR design offers better invariance and intuitiveness. The degradation caused by using a smaller VLM indicates that the capability of VLM is necessary to effectively understand instructions and visual inputs. We also conducted ablation studies on the methods employed in GEA training, as shown in the last three rows of Table 5. Each method was found to be critically important in enhancing the training effectiveness of the GEA.
@@ 5 Conclusion
In this paper, we introduce RoBridge, a novel hierarchical intelligent architecture designed to enhance robotic manipulation by bridging the gap between high-level cognitive planning and low-level physical execution. The architecture integrates a high-level cognitive planner, an invariant operable representation, and a guided embodied agent, demonstrating significant advancements in task generalization and execution robustness. Through extensive experiments, RoBridge has shown superior performance and strong zero-shot generalization capabilities in unknown environments and novel tasks.
Acknowledgements
This work is supported by Scientific Research Innovation Capability Support Project for Young Faculty under Grant No.ZYGXQNJSKYCXNLZCXM-I28, National Natural Science Foundation of China (NSFC) under Grants No.62476293 and 62272494, Shenzhen Science and Technology Program under Grant No.GJHZ20220913142600001, Nansha Key R&D Program under Grant No.2022ZD014, General Embodied AI Center of Sun Yat-sen University, and China Postdoctoral Science Foundation under Grant No.2025M771522.
@@ B.2 Training Details
Figure 5: RL training details.
In our experiments on MetaWorld, we conducted training for 1M steps, with failure data sampling occurring every 100k steps. We added the RL training curve. As shown in Figure 5, most tasks can be trained in 200k timesteps. The success rate of $\pi_{e}$ in the training scenario is 90.8%. During real-world experiments, we fine-tuned the model for 2k steps. We will resize the image to 168$\times$168 and feed it to GEA for faster speed. Training $\pi_{e}$ and $\pi_{g}$ on an A100 GPU took 25 and 30 GPU hours, respectively. Finetuning on real data took just 1 GPU hour. Inference needs 6 GB GPU memory (except GPT-4o). Only GEA and Track-Anything run every frame, taking 60-80ms. VLM and SAM+GroundingDINO only run in the first frame of each primitive action, with VLM at 0.3s and SAM+GroundingDINO at 1s per run.